OpenStudy (anonymous):

Prove the following identities: Can anyone can help me here..i will give medals and i will be your fan

OpenStudy (anonymous):

\[\frac{ \sin +\tan }{ \cot + \csc }=\sin \tan \]

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (cwrw238):

this is a tricky one - i'll have a look at it now

OpenStudy (anonymous):

ok... i have 3 more identities here if you dont mind can we solve that all?

OpenStudy (cwrw238):

I can solve this one - first convert everything to sin's and cos's - i'll write s for sin and c for cos

OpenStudy (cwrw238):

left side = (s + s/c )/ (cs + 1/s0 right side = s * (s/c) so lets try and simplify the left side and get the right

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

= sc + s c + 1 s(sc + s) ----- / ---- = ------- c s c^2 + c

OpenStudy (anonymous):

is it possible that we convert the 2 sides? or can we solve first in one side?

OpenStudy (cwrw238):

= s^2c + s^2 s^2(c + 1) --------- = -------- = s^2 / c = right side c^2 + c c(c + 10

OpenStudy (cwrw238):

so its just algebra

OpenStudy (cwrw238):

- wher you see 10 its a typo - should be 1)

OpenStudy (cwrw238):

i've used the identities tan = sin/cos , cot = cos/sin and csc = 1/sin

OpenStudy (cwrw238):

check out my third post - more typos it should read left side = (s + c/s) / ( c/s + 1/s)

OpenStudy (cwrw238):

- because tan = s/c , cot = c/s and csc = 1/s

OpenStudy (cwrw238):

shoot - whats wrong with me that (c + s/c) in the first bracket!!!

OpenStudy (anonymous):

how did you derive s(sc+s)/c^2+cos to sin^2/cos

OpenStudy (anonymous):

in the right side

OpenStudy (cwrw238):

ok - first expand the top part = s^2 c + s^2 and bottom part is c^2 + c to give s^2 c + s^2 --------- c^2 + c now factor top and bottom s^2( c + 1) --------- c(c + 1) the (c +1) will cancel out to give s^2/c which is the right hand side

OpenStudy (cwrw238):

when you have a mix of sec, cots tans or whatever its often best to convert to sines and cosines then aim to convert the more complicated side to get the simpler one

OpenStudy (cwrw238):

do you follow it ok?

OpenStudy (cwrw238):

its a bit long winded i know - there might be a shorter way to do it but i cant think of one.

OpenStudy (anonymous):

ok ok i get it

OpenStudy (anonymous):

can we solve another?

OpenStudy (cwrw238):

well i 've got 10 minutes only i'm afraid but ok

OpenStudy (anonymous):

1-2c^2/sc = tans-cot

OpenStudy (anonymous):

can we prove only the left side?

OpenStudy (cwrw238):

well we need to convert one side to the other

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you post it all at once :D

OpenStudy (cwrw238):

RHS = s/c - c/s = s^2 - c^2 / sc = cos2x / sinx cos x

OpenStudy (cwrw238):

sorry I cant at the moment because the 'post' box is running off the screen so i cant click on it

OpenStudy (cwrw238):

LHS = cos 2x / sc so both sides are the same

OpenStudy (cwrw238):

I've used the 2 identities sin^2 x - cos^2 x = cos2x and 1 - 2cos^2 x = cos 2x

OpenStudy (cwrw238):

gotta go right now